### v2

# SOFIA/HAWC+ FAR-INFRARED POLARIMETRIC LARGE-AREA CMZ EXPLORATION (FIREPLACE) SURVEY III: FULL SURVEY DATA SET

Abstract
We present the second data release (DR2) of the Far-Infrared Polarimetric Large-Area CMZ Exploration (FIREPLACE) survey. This survey utilized the Stratospheric Observatory for Infrared Astronomy (SOFIA) High-resolution Airborne Wideband Camera plus (HAWC+) instrument at 214 µm (E-band) to observe dust polarization throughout the Central Molecular Zone (CMZ) of the Milky Way. DR2 consists of observations that were obtained in 2022 covering the region of the CMZ extending roughly from the Brick to the Sgr C molecular clouds (corresponding to a roughly 1 0.75 region of the sky). We combine DR2 with the first FIREPLACE data release covering the Sgr B2 region to obtain full coverage of the CMZ (a 1.5 0.75 region of the sky). After applying total and polarized intensity significance cuts on the full FIREPLACE data set we obtain 65,000 Nyquist-sampled polarization pseudovectors. The distribution of polarization pseudovectors confirms a bimodal distribution in the CMZ magnetic field orientations, recovering field components that are oriented predominantly parallel or perpendicular to the Galactic plane. These magnetic field orientations indicate possible connections between the previously observed parallel and perpendicular distributions. We also inspect the magnetic fields toward a set of prominent CMZ molecular clouds (the Brick, Three Little Pigs, 50 km s, Circum-nuclear Disk, CO 0.02-0.02, 20 km s, and Sgr C), revealing spatially varying magnetic fields that generally trace the morphologies of the clouds. We find evidence that compression from stellar winds and shear from tidal forces are prominent mechanisms influencing the structure of the magnetic fields observed within the clouds.
1 INTRODUCTION
The Central Molecular Zone (CMZ) of the Galactic Center (GC) is an extreme region of the Milky Way containing elevated molecular densities (10 – 10 cm; e.g., Mills et al., 2018) and magnetic field strengths (100s of G – mG, Yusef-Zadeh & Morris, 1987a; Chuss et al., 2003a; Pillai et al., 2015; Mangilli et al., 2019) compared to what is observed in the Galactic disk. A 3-color view of this region probed by radio and infrared observations is shown in Figure 1, where prominent molecular clouds throughout the CMZ have been labeled. The extreme properties of the CMZ, compared to those observed in the Galactic Disk, and the fact that it is the closest Galactic nuclear region to Earth, makes it an important region to study to deepen our understanding of galactic nuclear regions more generally.
The star formation rate (SFR) is much lower in the CMZ than in the Galactic disk (e.g. Morris, 2023; Longmore et al., 2013). The low SFR is particularly unexpected given the amount of dense gas in the CMZ (e.g. 2-610 M, Barnes et al., 2017; Morris & Serabyn, 1996). There are multiple possible explanations for the low SFR in the CMZ, such as the strength and compressibility of turbulence in the region or that the CMZ could be in a period of inactivity between episodic bursts of star formation (Krumholz & Kruijssen, 2015). Another possibility could be the elevated magnetic field strengths observed in the region.
The role of the magnetic field in the CMZ has been studied using both radio and infrared observations. Radio observations reveal a population of synchrotron structures known as non-thermal filaments (NTFs) that are understood to trace ordered magnetic fields. These NTFs are predominantly oriented perpendicular to the Galactic plane (Yusef-Zadeh et al., 2022; Heywood et al., 2022; Morris, 1996). The general orientation of the NTF population has been interpreted to indicate a large-scale poloidal magnetic field that pervades the GC (Morris, 2006; Yusef-Zadeh et al., 2004).
The magnetic field derived from studies of dust polarization at sub-millimeter and infrared wavelengths, conversely, reveal a magnetic field that seemingly conflicts with the poloidal magnetic field traced by the NTFs. The magnetic field derived from the dust polarization observations traces the most extended axes of the molecular clouds within the CMZ – an orientation that is largely parallel to the Galactic plane and consistent with a toroidal field (Guan et al., 2021; Mangilli et al., 2019; Nishiyama et al., 2010; Novak et al., 2000). It is unclear how the seemingly distinct magnetic fields traced by the infrared and radio observations are related; however, observations at 90 and 150 GHz reveal regions where the magnetic field becomes poloidal rather than toroidal (Guan et al., 2021). This could either indicate connections between the fields traced by dust polarization and the NTFs or a spatially variable distribution comprised of multiple magnetic field components (Guan et al., 2021). The latter is likely given the sensitivity of these observations to both thermal dust and synchrotron emission.
The Far-Infrared Polarimetric Large Area CMZ Exploration (FIREPLACE) survey was proposed to address the question of how the magnetic field at infrared wavelengths varies throughout the CMZ by utilizing observations made by the High-resolution Airborne Wideband Camera plus (HAWC+) instrument on the Stratospheric Observatory for Infrared Astronomy (SOFIA). The first FIREPLACE data release (DR1) covered a region of the CMZ centered on Sgr B2 (Butterfield et al., 2023, hereafter reffered to as FIREPLACE I). The field of view of these first FIREPLACE observations is marked by the green dashed rectangle in Figure 1. The DR1 observations reveal a more spatially varying magnetic field at 19.6″ resolution than is seen in previous CMZ-wide surveys (1′, Guan et al., 2021; Mangilli et al., 2019). An in-depth study of the magnetic field system local to one of the molecular structures identified in DR1 indicates that the magnetic field probed by FIREPLACE traces smaller-scale molecular structures local to the CMZ than in previous CMZ-wide surveys (Butterfield et al., 2024, hereafter referred to as FIREPLACE II).
In this paper we present the second FIREPLACE data release (DR2), which is comprised of SOFIA/HAWC+ observations covering the CMZ from the Brick to Sgr C (a roughly 1 0.75 region of the sky). We then combine the DR1 and DR2 FIREPLACE data sets to obtain a polarimetric map with full coverage of the CMZ (a 1.5 0.75 region). In Section 2 we describe the data-analysis procedure applied to the DR2 FIREPLACE observations. In Section 3 we describe the data quality checks performed to verify the reduction procedure used to make maps from the raw observations obtained from SOFIA/HAWC+. The CMZ-wide FIREPLACE observations are then presented in Section 4, where the reduced DR1 and DR2 data sets are combined. In Section 5 we study the detailed polarimetric properties of prominent molecular clouds in our DR2 data set and we present our conclusions in Section 6.
2 OBSERVATIONS AND DATA REDUCTION
2.1 DR2 Observations: 2022 SOFIA Flights
FIREPLACE DR2 is an expansion of the FIREPLACE pilot program. The DR1 pilot program observed the eastern third of the CMZ (the greater Sgr B2 complex) during SOFIA flights F775 and F777 on August 31, 2021 and September 2 2021, respectively. This DR2 release incorporates observations obtained from several 2022 SOFIA flights: four Southern hemisphere Christchurch-based flights in June and July 2022 (flights F890, F891, F893, and F895) and three Northern hemisphere Palmdale-based flights in September 2022 (flights F916, F917, and F918).
The full FIREPLACE survey covers the central 1.5 0.75 of the CMZ. This is the largest polarimetric map made by SOFIA. Given both the large map size and the extended flux in this region at 214 µm, the standard Chop-Nod-Match (CNM Harper et al., 2018) observing strategy was not practical. Instead, FIREPLACE utilized the on-the-fly mapping (OTFMAP) mode. The OTFMAP mode involves rapid scanning of the array over the field to create time streams that are subsequently fit for correlated noise components and map values using the SOFIA data reduction pipeline, which is based on the CRUSH algorithm (Kovács, 2008). This is done for each of four angular positions of the half-wave plate (0, 22.5, 45, and 67.5), and the Stokes , , and parameters are fit to the resulting images.
Even with the OTFMAP mode strategy, the full extent of the CMZ is larger than what can be feasible observed with a single scan. Therefore we developed an observing plan utilizing a set of fields that fully cover the CMZ. The fields are reduced using the SOFIA OTFMAP reduction software as described above and then merged into a final map. This observation approach is novel, and there is significant correlated noise on short timescales. Therefore, care has to be taken in optimizing the observations and verifying the data set.
For FIREPLACE, each scan is designed to cross the Galactic plane at an angle of either 30 relative to Galactic North. For each scan, we follow the same parameters as FIREPLACE I – the boresight of the detector follows a Lissajous scan with amplitudes of 27′ 4.5′. The ratio of the frequencies between the parametric sinusoidal curves of the Lissajous pattern is set to , with the higher frequency corresponding to the long axis. The total time for each scan is 120 seconds, and the scan rate is 300″ s. The planned boresight path for each field of view is shown in Figure 2. The numbering in the figure marks the center of each field and the area covered by each boresight trajectory is indicated by the semi-transparent rectangles. Each field was observed using two different orientations that are oriented 30 with respect to Galactic North. The region targeted for polarimetry measurements is shown by the white contour (at a level of 7189 MJy sr at 250 µm). Each scan was designed with two considerations in mind. First, we required the scan to have sufficient extent so that on each scan the detector array crosses an effective “zero reference,” where the intensity is significantly lower than the target region. Second, we required significant overlap between adjacent and crossing fields for repeated measurements of any sky position in our target region. This latter point enables sufficient cross-linking between map pixels and detectors for the simultaneous fitting of all of the time streams for the map. Although the current version of the SOFIA pipeline does not allow for such simultaneous fitting, future data reduction efforts are being explored to take advantage of this aspect of the observation plan for increased fidelity, especially on larger spatial scales.
Observations in the region spanning fields 1 - 5 were taken in 2021 and comprise the DR1 FIREPLACE observations presented in FIREPLACE I. Note that the 2021 fields were implemented differently than those shown in Figure 2 due to setup issues during the 2021 observation runs (see Figure 2 of FIREPLACE I which shows the orientations of the DR1 observation fields). DR2 re-observed some of the fields covered in DR1 (fields 4 and 5), to increase the spatial overlap between the DR1 and DR2 observations, but primarily focuses on fields 6-18 to complete the map of the CMZ (Figure 2).
2.2 Data Reduction
To reduce the raw observations obtained from the SOFIA Infrared Science Archive at IPAC (IRSA) we use the HAWC+ Data Reduction Pipeline (DRP). We used DRP version 2.7.0, which is a Python-wrapper built on a Java version of the CRUSH reduction algorithm (Kovács, 2008). Since the latest version of of DRP (version 3.2.0) has not been verified for such large maps with extended emission we did not use this more recent version of DRP to reduce our HAWC+ observations. The use of DRP version 2.7.0 has the added benefit of maintaining consistency with DR1, which also used this DRP version.
Apart from FIREPLACE I, the OTFMAP polarimetry mode on SOFIA has previously been demonstrated only for sources that have spatial extents smaller than the instantaneous field-of-view of the HAWC+ detector (Lopez-Rodriguez et al., 2022). The abundance of diffuse, extended sources in the CMZ therefore presents a unique challenge, given that it is difficult for CRUSH to differentiate between these extended structures and correlated noise. In the reduction of our 2022 observations we closely follow and build on the methods from FIREPLACE I.
To reduce these observations using CRUSH we set the rounds parameter to “-rounds=85,” which increases the number of iterations CRUSH performs when numerically modeling the correlated noise. We also set the “-extended” option, which indicates that CRUSH needs to account for extended emission during data reduction to better preserve extended source emission when identifying correlated noise. Furthermore, we specified “-fixjumps” and “-downsample=1” to remove the effects of quantum flux jumps in the SQUID amplifiers and to prevent spatial averaging, respectively. These CRUSH settings match the parameters used for the FIREPLACE DR1 data release presented in FIREPLACE I.
Using these parameters in CRUSH, we reduced each field separately. For each field all scans were analyzed simultaneously where possible (8 scans for most fields accounting for 2 sets of 4 HWP angles). In some cases, an observation consisted of only one set of 4 scans. In these cases, 4 scans were included in the CRUSH step. Table A.2.1 in Appendix A.1 provides details on each of the scans observed and their mapping to the fields shown in Figure 2. Table A.2.1 also indicates how many scans were performed for each field.
Due to systematically high correlated noise that was evident in the post-reduction scans for the relevant fields, data from flight F890 were omitted from the final reduction. Fortunately, the regions observed during this flight were also observed during other flights, and so we still obtain full CMZ coverage. For completeness Table A.2.1 includes information on the observations conducted during flight F890, but we emphasize that the results presented in this work do not incorporate the observations obtained from that flight. The uncalibrated observations can be obtained from the SOFIA:IRSA data archive111https://irsa.ipac.caltech.edu/frontpage/ and can be found using the proposal ID of the survey (09_0054).
After reducing each field using CRUSH, the resulting total intensity emission of the fields was cross-correlated with Herschel 250 µm observations from Molinari et al. (2011) to identify pointing offsets in our fields. We found that the majority of fields corresponding to the 2022 observations had pointing offsets on the scales of 10″(corresponding to offsets 2 or more pixels in our map). These significant pointing offsets were corrected in all fields in which an offset of more than 5″ (angular size of 1 pixel in our map) was identified. Correcting the FIREPLACE pointings using the Herschel 250 µm observations completely corrects these offsets. Offsets smaller than 5″ are below the pixel size of our observations, so we did not implement offset corrections for these fields.
After applying the offset correction, the instrumental polarization is then removed from each field. The calibrated, offset-corrected, and instrumental polarization-corrected fields were then co-added into one combined data set using the DRP merge algorithm. Figure 3 shows the total intensity distribution for the merged 2022 observations with polarization pseudovectors shown as yellow lines. These polarization angles are calculated using
where and are defined following the IAU convention where North corresponds to an angle of 0 with increasing in the counterclockwise direction222The original reduction was done in equatorial coordinates. Thus “North” here refers to “equatorial North.” When switching to Galactic coordinates, the angle is rotated to be relative to “Galactic North.”.
The length of each pseudovector in Figure 3 is proportional to the debiased fractional polarization defined as
where the estimate of the uncertaintainty of the measured fractional polarization propagated from the fitted weights in the CRUSH step, and is the measured fractional polarization. Pseudovectors are only derived for lines of sight having total intensity above a significance threshold of , percentage polarizations less than 50% ( 50%), and significant polarized intensity (), which are quality cuts consistent with standard SOFIA polarimetry practice (Gordon et al., 2018).
3 DR2 Quality Checks
3.1 Testing Convergence of CRUSH Reduction
The large amount of extended emission present in our observations of the CMZ requires that we run CRUSH for more than the usual 15 rounds of fitting iterations. To confirm that the CRUSH fits are converging using this larger number of rounds, we conducted a second reduction of our HAWC+ observations that used the same CRUSH parameters (e.g. using the “extended” flag and setting “-downsample=1”) but changing the number of rounds to 75 instead of 85 (“-rounds=75” instead of “-rounds=85”), matching the reduction verification conducted in FIREPLACE I. We then created difference maps of Stokes , , , and of the polarization angle () between the 85 and 75 rounds reductions to assess whether the different numbers of iterations are converging to similar solutions. The difference maps for the merged 2022 flights are shown in Figure 4.
The top left panel of Figure 4 shows the difference in Stokes between the different number of rounds. These differences are generally 1000 MJy sr, and therefore the magnitudes of these differences are 1% of the total intensity of the significant emission obtained from our observations as can be seen from the colorscale of Figure 3. The magnitudes of the differences in the Stokes (upper right panel of Figure 4) and (lower left panel of Figure 4) distributions are an order of magnitude smaller than the differences observed for Stokes , and relative to the magnitudes of the Stokes and distributions, these differences are on the 1% level in regions of high linearly polarized emission. The polarization angle differences shown in the lower right panel of Figure 4 are generally 10. Since a 10 uncertainty in polarization angle corresponds to /, the variation resulting from the different numbers of rounds is smaller than the statistical noise resulting from the significance cuts employed in this paper. We therefore conclude that changing the number of iterations from 75 to 85 rounds is not significantly changing the results obtained from the CRUSH algorithm.
3.2 Variance Within DR2 Observations
We further test the convergence of the CRUSH fits by computing the variance for the percentage fractional and parameters across the entire CMZ. The percentage fractional Stokes parameters are defined such that: . We are able to compute variance maps of our observations because our observing strategy ensures that each pixel is sampled in multiple fields. The overlap between each field is such that an image made using every third field would have contiguous, non-redundant coverage over most of the CMZ. We therefore create sub-merges using every third field and for both of the 30 orientations, resulting in 6 sub-merges of the FIREPLACE observations that cover the extent of the CMZ observed in DR2. We compute the variance for each line-of-sight in our observations by treating the measurement from the full merged data set at that sky location as the “model” value, with each of our six data subsets being the different measurements of this “model” value.
The square-root of the variance values obtained for the percentage fractional and distributions are shown in the left and middle panels of Figure 5. The right panel of Figure 5 displays the hitmap indicating how many of the six subsets of observations overlap with each pixel evaluated in this procedure. The central region of this hitmap peaks at an overlap of six, but regions towards the edge of the map exhibit less overlap because of some fields that were removed from the final set of observations due to calibration issues. In addition, there is less overlap towards the center of the hitmap as a result of an unobserved field. The number of overlapping subsets also decreases towards the edge of the distribution because of how the fields are distributed and oriented on the sky (Figure 2).
We compute the variance only for the lines-of-sight that pass the significance thresholds discussed at the end of Section 2.2. In these regions of significant total and polarized intensity, the variance is generally 1% in polarization percentage. The variance is also generally quite flat, though there are some regions where the variance is . The small percentage variances observed indicate that the Stokes , , and emission is generally consistent across the different sub-portions of the combined 2022 data set; however, some areas of the map indicate the possible presence of systematic errors as can be seen where the and variance approaches 1%.
3.3 CNM Observing Mode Comparison
To further validate the quality of our reduction we follow the procedure employed in FIREPLACE I to compare our observations with archival CNM mode observations for 3 prominent molecular clouds covered by FIREPLACE DR2: the Brick, the Circum-nuclear Disk (CND), and Sgr C. Similar to the FIREPLACE data, these CNM mode data sets are also available on the SOFIA:IRSA data archive. The CNM mode data set for the Brick, consisting of 5 files, was obtained on flight 394 on May 12, 2017 by principle investigator T. Pillai (proposal ID: 0206). The CNM mode data for the CND and Sgr C regions were obtained on flight 397 on May 18, 2017 by principle investigator D. Chuss (proposal ID: 0018), consisting of 4 files for the CND and 3 files for Sgr C.
Figure 6 (left column) shows the inferred polarization pseudovectors of the CNM mode and OTFMAP mode data in (top) the Brick, (middle) the CND, and (bottom) Sgr C. The pseudovectors in these maps have been shifted slightly for ease of comparison. A polarimetric signal-to-noise cut for each data set, , and a total intensity signal-to-noise cut for the OTFMAP mode data set, , were applied, which are the same significance cuts employed in Section 2.2. The polarization angles of the CNM mode and OTFMAP mode observations generally agree for intensities greater than the red dashed vertical line in the right hand panels of Figure 6. Less agreement is seen around low intensity regions for all three clouds where the reference beam contamination for the CNM mode observations is potentially more significant (i.e. the spatial differences between the two observing methods is most impactful).
The right column of Figure 6 quantifies the comparison through calculation of the Pearson correlation between the normalized Stokes and parameters of the CNM mode and OTFMAP mode data as a function of a total intensity lower threshold. The Pearson- value is calculated only between points where the total intensity exceeds the intensity threshold, as determined from the OTFMAP mode total intensity map. The maximum intensity value was chosen to ensure that a minimum of 100 pixels (approximately 6 independent beams) are included in each correlation.
For intensities less than the red dashed vertical line in the right panels of Figure 6 the correlations are observed to drop. This drop is likely due to differences in spatial filtering between the OTFMAP and CNM mode observations, since spatial filtering is more significant in lower total intensity regions. At higher total intensities there are discrepancies between the Stokes parameters where one of the normalized Stokes parameters has a higher Pearson correlation than the other. For example, the correlations derived for the CND reveal a systematically higher correlation for normalized fractional Stokes than for normalized fractional Stokes . This discrepancy is caused by different signal-to-noise levels for the different normalized parameters where in the case for the CND there is generally much higher signal in than in .
In general, for all three clouds there is an increase in the Pearson correlation with increasing intensity. This trend is expected, since at higher total intensities contamination in the reference beam is less significant. The trend of increasing correlation with total intensity supports the validity of the data reduction pipeline employed for our FIREPLACE observations.
3.4 Comparison With Herschel 250 µm
As an additional check of the consistency of our new FIREPLACE observations, we plot the FIREPLACE DR2 Stokes (214 µm intensity) against the Herschel 250 µm intensity (Molinari et al., 2011) in Figure 7. The Pearson value is found to be 0.91, showing good correlation between the two data sets.
There are instances where the 214 µm intensity from FIREPLACE is 0 as can be seen in Figure 7. The negative emission indicates that the SOFIA/HAWC+ observations are insensitive to large-scale emission associated with very extended structures.
Though there is variation in dust temperature throughout the CMZ, we note that the slope of 0.91 found is consistent with dust having a temperature of 12 K with a spectral index of . This temperature is low for the molecular population in the GC, which exhibit an average temperature of 20 K (Molinari et al., 2011). This estimate of the slope is biased downward by the fact that the SOFIA/HAWC+ Stokes map is less sensitive to extended emission than the Herschel 250 µm map as described previously when discussing the negative 214 µm intensities observed in Figure 7.
Fitting only the positive values in Figure 7 would yield a larger slope for the fit to the correlation between FIREPLACE and Herschel. A higher slope would be consistent with dust possessing a higher temperature, and would therefore be more aligned with the average molecular temperature observed in the CMZ.
3.5 Validation Summary
As mentioned previously, the large area and prevalence of extended emission in the CMZ makes the polarimetric data reduction of our observations quite challenging. We have shown that the DRP reduction for this region is generally robust and the results are in good agreement with brighter molecular CMZ structures that have existing CNM mode observations at 214 µm. The reduced FIREPLACE data set presented here still likely contains residual systematics, and individual measurements in lower 214 µm intensity regions should be viewed with caution.
An additional challenge for the GC is the large amount of magnetized dust along the line-of-sight. By virtue of our reduction strategy, the spatial filtering of the FIREPLACE observations limits our sensitivity to features smaller than our beam size (19.6″). Future work that incorporates improved data reduction methods may yield a better understanding of the 3D magnetic field structure towards the GC.
4 The Polarimetric CMZ at 214 µm
Having verified our DR2 reduction we combine these DR2 data with the DR1 observations presented in FIREPLACE I for a full map of the CMZ. This combination is straightforward since the DR1 and DR2 fields were reduced using identical CRUSH parameters. We therefore co-added all of the individual DR1 and DR2 fields using the DRP merge algorithm. The combined fields consist of the DR2 fields presented previously in Table A.2.1 and the set of DR1 fields shown in Table A.2.2.
The upper panel of Figure 8 reveals the polarization pseudovectors (yellow dashes) obtained from our combined FIREPLACE DR1 and DR2 observations that cover the entire CMZ. Polarization orientations are only displayed for lines of sight that pass the selection criteria described in Section 2.2. Even with these significance cuts we obtain 65,000 Nyquist-sampled pseudovectors from our full 214 µm map of the CMZ.
Following the approach of FIREPLACE I, for the remainder of this paper we assume that the dust grain alignment mechanism responsible for the polarized emission is such that the angular momentum vector of the grains is predominantly parallel to the magnetic field direction and that the grains are spinning preferentially about their axis of greatest moment of inertia. This implies that the inferred direction of the magnetic field in the plane of the sky is perpendicular to the emergent polarization (or electric field) direction. This is consistent with -RAT alignment theory (Andersson et al., 2015; Lazarian & Hoang, 2007). We note that ultimately this interpretation of polarization will rely on a more complete theoretical understanding of the physics of grain alignment than currently exists. The lower panel of Figure 8 shows the derived magnetic field pseudovectors obtained by rotating the polarization angle orientations shown in the upper panel of Figure 8 by 90 under the assumption of -RAT alignment.
We display the Line Integral Contour (LIC) representation of these magnetic field pseudovectors in Figure 9, made using a kernel size of 1′. The LIC is a method of illustrating the polarization pseudovectors as streamlines (Cabral & Leedom, 1993). To do so, we filter the input distribution of polarization pseudovectors along local streamlines to produce an output image that represents the ‘flow’ direction of the polarization orientations. This visualization technique leads to a representation of the polarization that emphasizes the field geometry on the scales of molecular clouds in the CMZ.
The magnetic field orientation obtained from FIREPLACE is seen to generally trace the morphology of individual molecular clouds, as evident in the inset panels of Figure 9. The fact that the magnetic field orientation is impacted by the morphology of the CMZ molecular clouds indicates that the field derived in FIREPLACE is likely local to the CMZ rather than some other magnetic field component encountered along the line-of-sight.
A comparison of the FIREPLACE sample to previous infrared polarimetric data sets of individual molecular clouds within the CMZ made using the Kuiper Airborne Observatory (KAO), Caltech Submillimeter Observatory (CSO), and James Clerk Maxwell Telescope (JCMT) is shown in Figure 10. THE KAO and CSO pseudovectors are scaled like the FIREPLACE pseudovectors: proportional to fractional polarization. The JCMT pseudovectors, conversely, are not scaled to match their presentation in Lu et al. (2023). It is clear from Figure 10 that the FIREPLACE survey provides extensive polarimetric coverage of the CMZ, yielding a large sample of pseudovectors that will be used for statistical analysis. Figure 10 clearly reveals the sensitivity improvement the FIREPLACE observations provide in probing the magnetic field local to the CMZ molecular cloud population. In particular, the FIREPLACE observations yield sensitivity to fainter dust emission regions further from the Galactic plane and between the prominent molecular clouds in the CMZ. This increased sensitivity is a result of our observing strategy utilizing OTFMAP mode polarimetry and overlapping fields. The FIREPLACE magnetic field directions are in general agreement with those of previous infrared polarimetric observations of prominent CMZ molecular clouds, as can be seen in Figure 10; however, there are examples where the field derived from FIREPLACE is systematically different than previous observations. Comparisons between the pseudovectors derived from the FIREPLACE observations and previous individual cloud observations are discussed in more detail in Section 5.
4.1 The CMZ Magnetic Field on Multiple Spatial Scales
We first compare the magnetic field orientations obtained from FIREPLACE at 19.6″ resolution with the magnetic fields derived from ACTPol at 1′ and PILOT at 2′ resolution (Guan et al., 2021; Mangilli et al., 2019, respectively). This comparison is shown in Figure 11 where the magnetic fields are superposed on the 250 µm total intensity () CMZ emission from Herschel (Molinari et al., 2011). In general the magnetic fields from PILOT and ACTPol are smooth and parallel to one another, indicating that they are tracing a similar magnetic field component. The magnetic field traced by FIREPLACE is more spatially variant than either the PILOT or ACTPol distributions. The FIREPLACE magnetic field tends to trace the total intensity morphology of prominent CMZ molecular clouds as seen in high emissivity regions in Figure 11 and previously in Figure 9. In lower emissivity regions the magnetic field does not exhibit an apparent preferential orientation, ranging from being parallel to the Galactic plane to perpendicular to the Galactic plane.
Because of the sensitivity of ACTPol and PILOT to large-scale polarization relative to FIREPLACE, the magnetic fields derived from these surveys likely includes a magnetic field contribution originating from emission external to the CMZ. In other words, the increased spatial filtering and higher resolution of FIREPLACE relative to PILOT and ACTPol likely results in measurements of polarization local to the dust structures in the CMZ, whereas the PILOT and ACTPol measurements include contributions from a larger part of the central kpc of the Galaxy. Careful modeling of potential line-of-sight magnetic field components for these data sets will enable an analysis of the magnetic field configuration along the line-of-sight. For example, comparison of IR emission from stars on the near and far side of the CMZ indicates that this field component is connected to material that is within the central 12 kpc of the Galaxy (Nishiyama et al., 2009).
4.2 The Orientation of the Magnetic Field Throughout the CMZ
Building on the comparison of magnetic field scales presented in Section 4.1, we study the distribution of projected polarization pseudovectors to determine systematic trends in the polarization angle (and by extension the magnetic field) throughout the CMZ. The objective here is to quantify the extent to which the magnetic field derived from FIREPLACE connects to the large-scale magnetic field observed by PILOT and ACTPol. We plot the total intensity versus polarization angle in Figure 12, where the individual data points are colored based on the polarization percentage (/100) for each line-of-sight. The polarization angle distribution covers the full range of angles from -90 to +90. Furthermore, an inspection of the observed polarization percentage of the observations reveals that polarization levels 20% are rare and only present at 3,000 MJy sr. Only measurements have a polarization percentage 20%, comprising 1% of our entire data set. These high polarization measurements are likely a result of systematic effects or statistical fluctuations. The 3000 MJy sr regime corresponds to the peripheries of the molecular clouds in the CMZ, revealing that the polarization percentage is generally higher in regions of lower total emission. This emission regime is demarcated at its upper bound by the lower black horizontal dashed line shown in Figure 12.
At the highest total intensities (30,000 MJy sr) the number of polarization angles obtained is more sparse. This total intensity regime corresponds to the brightest molecular clouds in the CMZ, including Sgr B2, Sgr C, and the Brick. The fractional polarizations are generally 5% in this regime. This high total intensity regime is demarcated at its lower bound by the upper horizontal dashed line in Figure 12. The generally lower polarization percentages in these denser regions relative to the percentages in the fainter cloud peripheries agrees with previous polarimetric results for CMZ molecular clouds (e.g. Chuss et al. 2003a, b; Pillai et al. 2015; Lu et al. 2023; FIREPLACE I).
The large number of individual lines of sight in Figure 12 (65,000 Nyquist-sampled pseudovectors) makes it difficult to assess whether there are any preferential polarization angle orientations in the CMZ. To obtain better insight into CMZ-wide polarization angle trends we sort the individual polarization angles shown in Figure 12 into 10 bins. The histogram representation of the binned polarization angles is shown in Figure 13 (similar to what is done in FIREPLACE I). In these histograms, an angle of 0 indicates a magnetic field that is aligned with the Galactic plane, with an angle of 90 indicating a magnetic field that is oriented perpendicular to the Galactic plane.
The top row of Figure 13 reveals the polarization angle distribution for all polarization pseudovectors from the CMZ (corresponding to the entire polarization angle distribution shown in Figure 12). The remaining rows reveal the polarization angle distributions for the different total intensity regimes marked in Figure 12.
Column A illustrates the fractional polarization distribution for vectors of each intensity range. We observe a steeper slope for the higher total intensities (where 30,000 MJy sr). This steeper distribution was also observed in FIREPLACE I and could indicate increased depolarization towards the brighter (higher density) regions. This increased depolarization could be a result of decreased grain alignment efficiency in the interiors of molecular clouds (Santos et al., 2019), superpositions of multiple field components along the line-of-sight, or field tangling on scales smaller than the 19.6″ SOFIA/HAWC+ beam size of these observations (e.g. Fissel et al., 2016). Higher resolution observations are needed to determine which of these mechanisms is responsible for the increased depolarization observed in our results. We see an increasingly shallow slope as we move to fainter intensity regimes as marked in the third and fourth rows of Figure 13. This trend agrees with what was observed in DR1 (FIREPLACE I).
The histograms presented in columns B - D reveal preferential polarization angle orientations. The histogram corresponding to all of the polarization pseudovectors (top row) reveals a bimodal distribution in the polarization angle, where the histogram peaks at an angle of approximately -25 and 90. This is a similar bimodal distribution to the one observed in FIREPLACE I. We observe the same polarization angle enhancements in the middle and low emission regimes as seen in rows 3 and 4 of Figure 13. In the high regime presented in the second row of Figure 13 there is no 90 enhancement and there is instead an enhancement at 0.
The peak at in the brightest intensity regime that corresponds to the emission originating from the central, brightest and densest regions of the prominent CMZ molecular clouds. Since this distribution is connected to high emission () regions we find it unlikely that the orientation at is a magnetic field component external to the CMZ.
We also bin the 2022 polarization angle distribution shown in Figure 3 to assess whether the same polarization angle enhancements are recovered in this portion of the FIREPLACE observations. These histograms are presented in Appendix A.2 and shown in Figure A.1. We find a similar bimodal polarization angle enhancement at angles of -20 and -80 in all 2022 pseudovectors and in the middle and low emission regimes. These are similar bimodal enhancements seen in the full FIREPLACE data set in these intensity regimes. However, in the high emission regime we do not detect the 0 enhancement and instead observe an enhancement at -60. Therefore, the 0 component seen in the emission regime of the full data set can be associated with Sgr B2. Since this high intensity regime corresponds to the cores of the brightest and most dense molecular clouds, the different enhancements observed in this regime indicate that the magnetic field is preferentially aligned along the morphology of these clouds, further corroborating that the magnetic field in this regime is local to the CMZ.
To further quantify the bimodal distribution from the full FIREPLACE data set, we fit the histogram of polarization pseudovectors obtained for the full FIREPLACE data set with a two-Gaussian model. To perform this fitting, we assume the uncertainties on the counts in each histogram bin is the square-root of the counts in that bin (ranging from 44100 counts). These fits are shown in Figure 14, where the polarization angles derived from FIREPLACE have been sorted into 5 bins for the full set of polarization angles and the middle intensity regime. The blue line is the best-fit result obtained from fitting a two-Gaussian model to the histogram of the polarization angle data defined as:
where is the offset above zero for the Gaussian components (in counts), and are the amplitudes of each Gaussian (in counts), is the set of polarization angles from -90 to +90 that comprise the histogram, and are the centroids of each Gaussian (in ), and and are the standard deviations of each Gaussian (in ). This results in a seven-parameter model, with all eight parameters allowed to vary to determine the best-fit to the underlying histogram. The resulting best fit parameters for the full polarization vector distribution and the middle intensity regime are shown in Table 1.
The peaks of the modelled Gaussian components are located at polarization angles of and . The residuals are shown below the model fit, and range from a few hundred to 1000 counts. These residuals are fairly large compared to the uncertainties in the histogram bins, but this two-Gaussian model fit is simple, and is likely not fully characterizing the structure of the histogram.
The histogram and resulting best-fit model for the middle intensity regime corresponding to the third row of Figure 13 is shown in the right panel of Figure 14. The fit returns similar parameters that are consistent within the error bars of the two measurements with peaks located at and . As with the full polarization angle pseudovector, the residuals corresponding to the best-fit model are implying that the model is under-fitting the structure of the histogram. More detailed characterization of these histograms would require a model consisting of additional model components. Theoretical modeling is needed to estimate possible additional model components.
The polarization angle peak from the model fit at approximately corresponds to a magnetic field orientation that is aligned with the CMZ-wide magnetic field observed by the ACTPol and PILOT programs (Guan et al., 2021; Mangilli et al., 2019). As discussed this field is connected to material in the inner 12 kpc of the Galaxy (Nishiyama et al., 2009). This enhancement could therefore be tracing the orientation of the CMZ molecular clouds or it could be a magnetic field component external to the CMZ. The Gaussian component centered at approximately relates, conversely, to a vertical magnetic field and provides evidence of a connection between the magnetic fields in the dust to those traced by the radio NTF population. We do not fit a two-Gaussian model to the high-intensity regime because of the lower number of polarization angles per histogram bin in this regime as can be seen in Figure 13 – 100 compared to the full distribution with 1000.
The high-total intensity regime cutoff of 310 MJy sr is lower than was used in FIREPLACE I, where a threshold of 510 MJy sr was used to isolate the magnetic field orientation obtained from only the brightest regions of Sgr B2. As a further verification of the combined FIREPLACE data set we analyze the histogram of polarization angles obtained from using the DR1 upper total intensity regime of 510 MJy sr to isolate the polarization angles from Sgr B2. We would expect to obtain the same distribution as shown in FIREPLACE I. We find the same number of polarization pseudovectors to what was found in DR1 (359) corresponding to this regime where 510. We find essentially the same histogram distribution as reported in Figure 8 of FIREPLACE I. There is some variation where certain 10 bins are one count higher or lower in our combined observations compared to the histogram presented in FIREPLACE I. Such slight differences in individual histogram bins are likely a result of slight rounding differences where polarization angles near the 10 bin edges are placed in different bins.
The orientation of the magnetic field enhancements help to clarify a picture of the magnetic fields across the central degree of the Galaxy. In all histograms, we see excess measurements at an angle consistent with that of the large-scale component seen in the PILOT and ACTPol data (the -20 enhancement discussed for Figures 13, 14, A.1). In the high intensity regime (the yellow histogram of Figure 13), the polarization angle enhancements observed at -20 splits into two sub-components: one at -20 and one at 0. However, in the 2022 high intensity portion shown in Figure A.1 the enhancements are at -60 and -20.
The 0enhancement is only observed in the brightest portion of Sgr B2 is almost certainly associated with the core of the Sgr B2 region, since we do not recover this component in the 2022 observations (Figure A.1). The -20 component is detected in all our intensity regimes. In the high intensity regime the -20 enhancement is likely tracing a local CMZ field. For the lower intensity regimes, it could be at least partially an external magnetic field component given the large spatial extent over which this field is coherent at arcminute resolutions. The enhancement at 90 is only detected in the middle and low intensity regimes. The detection of this enhancement in only these regimes is consistent with the picture of a connection between the vertical field traced by the NTFs and a horizontal field that is a result of the field being sheared in denser, kinematically-dominated regions.
Previously, it has been posited that an initially poloidal field could be sheared into a horizontal configuration in denser molecular regions (Morris & Serabyn, 1996). Our findings here support early submillimeter data (Novak et al., 2003; Chuss et al., 2003a) and near-infrared polarimetry data (Nishiyama et al., 2010) that have suggested a transition occurs from a sheared field in dense regions to a vertical field in regions having lower kinetic energy density. The two scenarios that have been considered are the above-mentioned evolution from an initially poloidal field and the creation of a vertical field from the sheared field via winds and outflows.
In producing these histograms, we have utilized polarimetry data across the range of intensities down to 1000 MJy sr. If systematic errors remain after our data cuts, they are unlikely to produce the histogram peaks observed. The scan angles () are sufficiently different from the maxima of the histograms that a systematic scan-synchronous effect is unlikely to produce the results seen here.
| Parameter | All Polarization Angles | Middle Intensity Regime |
| (counts) | 3810 | 1850 |
| (counts) | 5780 | 4250 |
| (counts) | 6720 | 5450 |
| () | -82.1 | -84.3 |
| () | -18.6 | -19.6 |
| () | 17.7 | 16.4 |
| () | 22.4 | 24.4 |
5 THE MAGNETIC FIELDS OF PROMINENT CMZ MOLECULAR CLOUDS
We have argued the FIREPLACE results predominantly reveal the magnetic field local to the CMZ, especially in bright emission regions. In this section, we compare the FIREPLACE 214 µm results with previous polarimetric observations of individual molecular clouds. Furthermore, the increased sensitivity of the FIREPLACE survey enables a determination of how the magnetic field local to the CMZ varies in the peripheries of these clouds than previous studies. Here we present an overview of the magnetic field configurations obtained for a set of prominent CMZ molecular clouds, as indicated in Figure 15.
In the following discussion Figures 16 - 20 are used to motivate the discussion of the magnetic fields local to the distinct clouds. These figures all show how the FIREPLACE magnetic field pseudovectors compare to previous polarimetric observations of these clouds. In general, our new 214 µm observations find a magnetic field that is consistent with these previous studies. Cases where this is not the case are discussed in the relevant subsections.
5.1 M0.253+0.016 (The Brick)
M0.253+0.016, known colloquially as “The Brick,” lies to the Galactic West of the Dust Ridge and is near the arched filaments that are proximal to the Radio Arc NTFs (Lang et al., 1999; Yusef-Zadeh & Morris, 1987b). The location of the Brick in the GC can be observed in Figure 15. A zoom-in view of the Brick’s total intensity emission obtained from Herschel at 250 µm is shown in Figure 16. Overlaid on this total intensity distribution are the magnetic field pseudovectors derived from our FIREPLACE observations (blue line segments), the magnetic field pseudovectors derived from 350 µm CSO observations (red line segments, Dotson et al., 2010; Pillai et al., 2015), and the magnetic field pseudovectors derived from 850 µm JCMT observations (cyan line segments, Lu et al., 2023). We obtain 50 magnetic field pseudovectors corresponding to the Brick, extending into fainter emission regions than the CSO or JCMT observations. The FIREPLACE pseudovectors are observed to trace the total intensity morphology of the Brick, generally following the extended curvature of the cloud. Our FIREPLACE magnetic field orientations also agree with those derived from the HAWC+ CNM mode observations of this cloud as seen in Figure 6.
The ordered magnetic field obtained from the CSO observations led Pillai et al. (2015) to postulate that the magnetic pressure exceeds the turbulent pressure in this cloud. If true, this indicates that the magnetic field is more dynamically important than turbulence and plays a more significant role in shaping the evolution of the Brick. The similarly ordered magnetic field derived from our FIREPLACE observations support this possibility, and we furthermore observe that the ordered field extends into fainter regions of the Brick than was measured by Pillai et al. (2015). Additionally, HCO+ absorption filaments observed within the Brick also appear to trace the magnetic field orientation within the cloud (Bally et al., 2014), which indicates that the density structures within the cloud could be shaped by the internal magnetic field geometry.
Pillai et al. (2015) argue that the curvature of the magnetic field seen towards the Brick implies the presence of a shock. A recent study of the kinematics and structure of the Brick reveals that the cloud is likely a conglomeration of distinct velocity components along the line-of-sight (Henshaw et al., 2019, 2022), which could result from the Brick consisting of distinct molecular structures that, although superimposed along the line-of-sight, are at different locations in the gravitational potential of the CMZ.
There is also evidence of feedback in the Brick given the presence of an expanding bubble likely driven by the stellar wind of a high mass star (Henshaw et al., 2022). The extent of this wind-blown bubble is marked with a red dashed circle in Figure 16. Feedback shocks resulting from the expansion of this wind-blown bubble could generate the shocked gas and explain the ordered and curved magnetic field of the Brick. In fact, inspection of the magnetic field pseudovectors that coincide with the periphery of the bubble reveals that they largely trace the periphery of the bubble, particularly in high emission regions coincident with the core of the Brick molecular cloud.
The magnetic field external to the periphery of the wind-blown bubble parallels the largely azimuthal magnetic field geometry seen for the M0.8-0.2 Ring near Sgr B2 that is studied in FIREPLACE II. The M0.8-0.2 ring is likely a supernova remnant, as shown from the multi-wavelength results discussed in FIREPLACE II. They report that the magnetic field in M0.8-0.2 becomes enhanced as a result of compression of the magnetic field due to the expansion of the shell. Similar physics could be occurring with the wind-blown bubble within the Brick.
The polarization fractions from the Brick are observed to increase towards the periphery of the cloud (5%), whereas the fractional polarizations in the core of the cloud are generally lower (1%). Increased polarization fraction towards the periphery of the Brick is consistent with the CMZ-wide trend of increased depolarization with higher emission as shown in Figure 13.
5.2 M0.084-0.081 (The Three Little Pigs)
M0.145-0.086 (Straw), M0.106-0.082 (Sticks), and M0.068-0.075 (Stone) together comprise the cloud complex known as the “Three Little Pigs” (TLP, Battersby et al., 2020). These clouds are labelled and shown in Figure 17. This complex is oriented nearly parallel to the Galactic plane. The clouds all have similar velocities which indicates they are associated with each other and not simply close in projection. The TLP is located 10′ in projection from Sgr A as can be seen in Figure 15.
Previous polarimetric observations of the TLP clouds yielded preliminary evidence of a bending magnetic field within the Stone cloud in the TLP and on the periphery of this cloud closest to Sgr A (Chuss et al., 2003a, b; Lu et al., 2023). the Straw and the Sticks clouds, conversely, were observed to have magnetic fields that are oriented largely perpendicular to the Galactic plane and do not exhibit as much warping as seen in the Stone cloud. This result is also observed in the arcminute-resolution ACTPol results presented in Guan et al. (2021). These previous polarimetric observations were either only sensitive to the brightest regions of the clouds in the TLP or were limited by lower angular resolutions, and so it was unclear from those observations how the field behaves between the clouds of the TLP and the outer periphery of the TLP complex.
Our new FIREPLACE observations provide over 100 magnetic field pseudovectors corresponding to the TLP that uncover the magnetic field between the TLP clouds and in the lower emission regions at the periphery of the cloud complex. The FIREPLACE 214 µm dust polarimetry observations recover magnetic field orientations in the cores of the Straw, the Sticks, and the Stone clouds similar to those derived from the CSO and JCMT observations.
The magnetic field surrounding the Straw cloud derived from FIREPLACE is perpendicular to the Galactic plane at lower Galactic latitudes (-0.1 and below in the figure) whereas the magnetic field above the cloud is angled almost parallel to the Galactic plane. The field above and below the Sticks cloud is oriented perpendicular to the Galactic plane and only becomes warped for lines of sight where the emission exceeds MJy sr. The field surrounding the Stone cloud exhibits the same curve as is observed in the core of the cloud from the previous CSO and JCMT observations.
The systematic differences in magnetic field orientation seen above and below the Straw cloud coincide with the low-longitude outer extent of the M0.20-0.030 shell likely originating from the Quintuplet cluster (Butterfield et al., 2022, 2018). Butterfield et al. (2018) argued that the TLP clouds could be interacting with the M0.20-0.033 shell. The magnetic field pseudovectors from FIREPLACE do appear to be impacted by the presence of the shell: changing from a systematically perpendicular (with respect to the Galactic plane) magnetic field orientation outside the shell to parallel or rotated within the shell.
The impact of the shell on the magnetic field pseudovectors is also observed in higher emission regions of the Straw cloud where the core of the cloud coincides with the outer extent of the shell. This result corroborates the argument of Butterfield et al. (2018) of an interaction between the shell and the TLP complex; indicating that at least the Straw cloud is likely impacted by the presence of the shell. A similar argument is made in Lu et al. (2023) based on their JCMT observations of this cloud.
Neither the Sticks nor the Stone clouds show any evidence of interaction with the M0.20-0.030 shell. However, the field coinciding with and surrounding the Stone cloud is noticeably more warped than that of the Sticks cloud. The Stone cloud in particular is too far in projection from the M0.20-0.033 shell for the curvature of the magnetic field in that region to be explained by interaction with that shell. Instead, the curvature could be the deformation of an initially vertical magnetic field induced by the orbital motion of the TLP cloud complex as it moves through the field. Additional analysis is needed to determine the connection between the properties of the TLP clouds and the larger field structure.
An explanation for the differences in the magnetic field geometries observed between the Straw, the Sticks, and the Stone cloud could be differences in mass between the different clouds. There is a distinct brightness gradient in the emission from the clouds in the TLP (where the brightness here is proportional to the column density), with the Straw cloud being the faintest and M0.068-0.75 being the brightest. The differences in mass for the clouds in the TLP could indicate changes in the magnetic critical parameter which quantifies the relative importance of the magnetic field and self-gravity within these clouds (Guerra et al., 2023).
A difference in the critical parameter between the clouds could indicate changes in the tendency of the clouds to contract and form stars (Nakano & Nakamura, 1978). This possibility is consistent with recent star formation rate (SFR) estimates obtained by the CMZoom survey that found an SFR of 6.43.8 M yr for the Stone cloud compared to 0.40.3 M yr for the Straw cloud (Hatchfield et al., 2023). The higher SFR of the Stone cloud increases the likelihood of there being outflowing winds within this cloud, similar to the bubble observed by Henshaw et al. (2019) within the Brick. Follow-up work is necessary to determine whether it is the orbital motion of the TLP or differences in the presence of stellar winds between the clouds that best explains the differences in the magnetic field seen for the Sticks and the Stone cloud.
As with the Brick we observe increased polarization fractions towards the periphery of the TLP cloud complex. The highest polarization fractions observed for this cloud are 5%, with regions in the center of the cloud having generally lower polarization fractions (1%) than what is observed in the fainter peripheries of the cloud complex.
5.3 Clouds Surrounding Sgr A
There are multiple clouds that are within 5′ in projection from Sgr A. The prominent clouds that are within this region and shown in Figure 15 are M0.001-0.058 (50 km s cloud), M359.948-0.052 (Circum-nuclear Disk), CO 0.02-0.02 (as labeled in Lu et al. 2023), and M359.889-0.093 (20 km s cloud). We analyze these clouds as a group and they are presented in Figures 18 and 19.
5.3.1 M0.001-0.058 (50 km s Cloud)
The 50 km s cloud is located to the Galactic Southeast of Sgr A and is contained within the greater Sgr A complex. The location of the 50 km s cloud is shown in Figure 15.
We obtain 50 magnetic field pseudovectors from our 214 µm observations and are largely oriented perpendicular to the Galactic plane. There are deviations from this perpendicular field orientation. For example, there is a dense dust ridge that lies along the Eastern periphery of the Sgr A East supernova remnant (SNR Serabyn et al., 1992) at = 359.967, = -0.067. The magnetic field coincident with this dust ridge traces the orientation of this ridge, which is rotated by 45 relative to the Galactic plane. This dust ridge in the 50 km s cloud is thought to be compressed by a SNR (Ehlerová et al., 2022; Novak et al., 2000; Serabyn et al., 1992), which can explain why the magnetic field is oriented along the ridge.
The magnetic field observed in the 50 km s cloud changes direction at Southern Galactic latitudes where the intensity decreases and the previous CSO and JCMT observations lack sensitivity. This is particularly noticeable at = 359.97, = -0.117 where the field subtends an angle of 60 with respect to the Galactic plane. The fractional polarizations are also quite large in this region (generally 5%). The magnetic field orientation also systematically changes from being generally oriented perpendicular to the Galactic plane to the East of the 50 km s cloud to being parallel to the Galactic plane to the West of the 50 km s cloud.
5.3.2 M359.948-0.052 (Circum-nuclear Disk)
The Circum-nuclear Disk (CND) is a set of clouds located within 5 pc of Sgr A. The CND has an inclination of 70 relative to the plane of the sky and an inner radius of pc (Guerra et al., 2023; Lau et al., 2013). The CND is labeled in Figure 18.
Continuum Stokes imaging with HAWC+ at 19.7, 31.5, 37.1, and 53 µm show a very clear set of clouds that are gravitationally related (either orbiting or infalling) to Sgr A* (Lau et al., 2013; Guerra et al., 2023) and are experiencing gravitational shear. At far-infrared wavelengths Molinari et al. (2011), the structure of the CND is only faintly apparent, with the emission being dominated by more extended, cooler CMZ dust, which is located external to the central 5 pc. The 214 µm Stokes maps are consistent with these images.
In polarimetry, the CND has been measured at 53 µm at 5″ resolution. These pseudovectors show a field that is highly aligned with the dust streamer morphology (Guerra et al., 2023) who argue for a steady-state shear flow leading to a magnetic field that traces the morphology of the CND arms. They estimate a milligauss magnetic field strength for the fields in the CND using a modified version of the Davis-Chandrasekhar-Fermi technique (Davis, 1951; Chandrasekhar & Fermi, 1953) that takes shear into account.
The FIREPLACE dataset contains 50 Nyquist-sampled magnetic field pseudovectors in the vicinity of the CND. In contrast to the shorter wavelength observations, the inferred magnetic field is quite ordered and is largely oriented parallel to the projected long axis of the CND, which itself is oriented at an angle of about 20 degrees with respect to the Galactic plane.
The FIREPLACE 214 µm data is consistent with the geometry measured at 350 µm at 20″ resolution (Novak et al., 2000). Parts of the CND were included in the SCUBA-2/Pol-2 map at 850 µm and 20″ resolution (Lu et al., 2023). There is general agreement with these vectors in the Northeastern part of this region; however, to the south, the 850 µm vectors do not agree with the shorter wavelengths. This could be an intrinsic variation in the source polarization at these different wavelengths or it could be a result of differences in spatial filtering from the different observing modes (OTFMAP map for this study compared to CNM mode for the CSO and Daisy mode for the JCMT (Lu et al., 2023)). However, our FIREPLACE results agree with the HAWC+ CNM mode observations of this cloud as seen in Figure 6. Therefore the difference is likely a result of changes in the polarization over the different wavelengths. Previous 850 µm polarization measurements have been made of the CND region (Hsieh et al., 2018), in which the authors claimed to detect a sheared geometry of the field.
The differences in morphology and polarization orientation from 53 to 214 µm (and longer wavelengths) are most likely due to two cloud components superposed along the line-of-sight the warm dust local to the streamers of the CND and the cooler dust component in the wider CMZ. The former contains magnetic fields that dynamically connected to the sheared flow of the inner 5pc; the latter are likely shaped by the dynamics of the greater CMZ.
5.3.3 CO 0.02-0.02
Cloud CO 0.02-0.02 is only 4′ away from the CND in projection to the Northeast of the CND as can be seen in Figure 18. The FIREPLACE observations reveal 20 magnetic field pseudovectors coinciding with the CO 0.02-0.02 cloud that are generally oriented perpendicular to the Galactic plane; however, this orientation does not agree with what is observed in previous CSO and JCMT observations of this cloud. These pseudovectors differ by a 30 angle with respect to the FIREPLACE pseudovectors. In addition, there are instances where the CSO and JCMT observations disagree, which could indicate complexity in this region of the cloud.
The disagreement in magnetic field orientation could be a result of different spatial filtering between our FIREPLACE observations and that of the CSO and JCMT observations. In addition, we note that the CO 0.02-0.02 cloud falls within the region of elevated and variance observed in Figure 5 from our FIREPLACE observations. The angle differences observed for this cloud could therefore indicate remaining systematic uncertainties in our FIREPLACE observations. This explanation is further justified by the fact that the CSO and JCMT observations are in general agreement over most of this cloud.
5.3.4 M359.889-0.093 (20 km s Cloud)
M359.889-0.093, known colloquially as the 20 km s cloud, is located 10′ in projection from Sgr A to the Southwest as shown in Figure 15. A zoom-in view of the 20 km s cloud is shown in Figure 19. The orange arrow in Figure 19 indicates the direction to Sgr A.
Our SOFIA/HAWC+ observations yield 100 distinct pseudovectors that span the full extent of the 20 km s cloud. The field within the cloud is predominantly oriented parallel to the Galactic plane. The field within the cloud is seemingly preserved above and below the cloud in Galactic latitude, being also largely oriented parallel to the Galactic plane.
The core of the 20 km s cloud is comprised of a set of filamentary structures in the dust continuum that appear to merge at =359.867, =-0.083. At this location we observe low fractional polarization values (1%), suggesting that depolarization in this region may be present. This depolarization could be caused by the overlapping of field directions along our line-of-sight from these multiple filamentary structures. Alternative explanations for depolarization in this region could be caused by loss of grain alignment in the densest regions of the cloud or multiple field components within the size of our beam. Depolarization in these denser regions has been observed in several of CMZ clouds including Sgr B2 and Dust Ridge cloud E/F (FIREPLACE I). Previous studies using combined SMA and VLA observations find that the star formation of the 20 km s cloud is likely in an early evolutionary stage (Lu et al., 2015). Sgr B2 and the Dust Ridge cloud E/F, which also show depolarization in their cores show signs of early and on-going star formation - indicating that the 20 km s cloud may be at a similar evolutionary state.
5.4 M359.484-0.132 (Sgr C)
Sgr C is located 0.5 in projection from the 20 km s cloud and is on the Western periphery of the CMZ as shown in Figure 15. The cloud manifests a mix of diffuse and compact emission coupled with a broad range of emissivities as shown in Figure 20. This cloud was recently observed polarimetrically using the JCMT (Lu et al., 2023) and the CNM mode observations made using SOFIA/HAWC+ that are shown in Figure 6. Prior to these observations Sgr C had not been studied polarimetrically at a resolution less than 1′. The FIREPLACE observations of Sgr C yield 100 magnetic field pseudovectors within the periphery of the cloud. These observations largely agree with the orientations derived from the SOFIA/HAWC+ CNM observations of this cloud as seen in Figure 6.
The emission from Sgr C obtained by Herschel is generally diffuse, but there is a region of compact, high intensity emission at = 359.433, = -0.1. There is a ridge or tail of emission emanating from this compact emission towards the Southeast, which we hereafter refer to as the “head-tail structure” (labeled in Figure 20). The magnetic field derived from our 214 µm observations follows the morphology of the head-tail structure and is largely consistent with the magnetic field orientation derived from the 850 µm JCMT observations of Lu et al. (2023). Lu et al. (2023) argue that the magnetic field orientation in the head-tail structure could result from interaction of the Sgr C molecular cloud with a nearby prominent H II region, the extent of which has been marked with an orange circle in Figure 20 (Hankins et al., 2020; Lang et al., 2010). Evidence of this interaction between Sgr C and the H II region has previously been presented in (e.g. Lu et al., 2019). The similar magnetic field orientations derived from FIREPLACE bolster the argument that winds from the prominent H II region create molecular pillars leading to a compressed magnetic field that follows the orientation of the head-tail structure. Magnetohydrodynamic simulations of expanding H II regions reveal similar molecular and magnetic field morphologies to what is seen in the head-tail structure of Sgr C (Arthur et al., 2011; Mackey & Lim, 2011).
Our results corroborate the simulation results obtained from these magnetohydrodynamic models of expanding H II regions. Furthermore, the field geometry seen in the high emission is preserved in lower intensities as measured by our FIREPLACE observations. Arthur et al. (2011) and Mackey & Lim (2011) argue that the magnetic field is dragged along the head-tail structure by the expanding wind from the H II region.
The magnetic field revealed by our SOFIA/HAWC+ observations Southeast of the head-tail structure at = 359.467, = -0.15 is largely oriented perpendicular to the Galactic plane. The JCMT pseudovectors in this Southern portion of Sgr C generally do not agree with those obtained in the FIREPLACE survey and are more spatially varying. To help understand the difference in polarization angle between the FIREPLACE and JCMT observations, we note that the significance of the JCMT polarized intensity in this region is reported as , whereas the significance of the JCMT polarized intensity in the head-tail structure in Sgr C is (Lu et al., 2023). is equivalent to a 10 uncertainty in the polarization angle, and the JCMT orientations disagree with those obtained from FIREPLACE by 70-90 in this Southern region.
To the Southwest of the head-tail structure our SOFIA/HAWC+ observations reveal a curved magnetic field structure at = 359.417, = -0.167. There are also JCMT pseudovectors at this location which are again highly spatial variable. The curvature in this region is similar to the field structures seen for the Brick and the TLP clouds that were attributed to out-flowing winds. Furthermore, a compact source from the ATLASGAL survey (AGAL359.474-0.152; Contreras et al., 2013), whose location is marked with a red ellipse in Figure 20, has an integrated flux of 168 Jy and a total gas mass of 5 10 M (Kendrew et al., 2013). These properties make AGAL359.474-0.152 favorable for forming massive stars (Lada et al., 2012).
Therefore, out-flowing winds from star formation in Sgr C, likely originating from AGAL359.474-0.152 is a reasonable origin for the structure of the magnetic field in this region of Sgr C, especially in light of the ample evidence of star formation in Sgr C (e.g. Lu et al., 2019; Yusef-Zadeh et al., 2009). However, there are alternative mechanisms that could result in the magnetic field distribution seen in this region, such as shear from orbital motion of the structure.
As with the previously discussed clouds, the fractional polarizations derived from our 214 µm SOFIA/HAWC+ observations increase towards the periphery of Sgr C in lower emission regions. In the region of diffuse emission to the South of the head-tail structure fractional polarizations range from 4 – 5%. In the head-tail structure, conversely, the fractional polarizations are 1%, similar to the peak emission regions of other clouds.
6 CONCLUSIONS
In this work the full FIREPLACE survey covering the entire CMZ polarimetrically at 214 µm is presented. First, the reduction method performed on the observations obtained in 2022 is detailed, comprising the second FIREPLACE data release (DR2) and observing the region of the CMZ extending from the Brick to Sgr C (corresponding to a 1 0.75 region of the sky). The reduced DR2 observations are then combined with the DR1 observations presented in FIREPLACE I to obtain a map of the entire CMZ polarimetrically at 214 µm (a 1.5 0.75 region of the sky). CMZ-wide trends of the magnetic field are then inferred and analyzed, with the fields local to a set of prominent molecular clouds inspected in more detail. The following summarizes the results from the inspection of the full FIREPLACE survey:
Our reduction method, using the SOFIA DRP software, results in a robust data set that is well verified and consistent with other SOFIA/HAWC+ CMZ polarimetric observations made using the CNM observing mode. The combination of the convergence, variance, and CNM mode comparison presented in Section 3 convincingly shows that our reduction is reliably removing the correlated noise while accurately recovering the source emission.
The sensitivity of our HAWC+ observations makes it possible to derive the polarization angle and magnetic field orientations in fainter molecular structures in the CMZ at a resolution of 19.6″ than ever before. The distribution of 65,000 Nyquist-sampled magnetic field pseudovectors from the full FIREPLACE survey is shown in Figure 10.
A bimodal distribution in the CMZ-wide magnetic field orientations is identified from the full FIREPLACE survey. This bimodal distribution shows enhancements in directions parallel (approximately -18) and perpendicular (approximately -86) to the Galactic plane. This is a similar bimodal distribution to the one recovered in FIREPLACE I for the Sgr B2 region. The component that is parallel to the Galactic plane could either indicate a toroidal magnetic field component in the CMZ (Guan et al., 2021; Mangilli et al., 2019) or a component external to the CMZ. The bimodal component that is perpendicular to the Galactic plane could be related to the poloidal magnetic field distribution in the GC (e.g. Morris, 2006), possibly revealing connections between the magnetic fields traced by infrared dust polarization and radio synchrotron emission.
The distributions of the magnetic fields of a set of prominent molecular clouds in the CMZ are also analyzed: the Brick, Three Little Pigs (TLP) cloud complex, 50 km s cloud, CND, CO 0.02-0.02, 20 km s cloud, and Sgr C. The magnetic field derived from FIREPLACE generally traces the morphology of the molecular clouds. In addition, there is evidence that expanding shock fronts from SNRs, advancing ionization fronts from H II regions, and tidal forces from orbital motion might cause compression and shearing leading to the organization of the magnetic fields in the Brick, TLP cloud complex, 50 km s cloud, and Sgr C. However, not all of the prominent clouds studied in this work exhibit signatures of compression or winds (the 20 km s cloud, for example). These results nonetheless indicate that the compression from winds and shear from tidal forces are important mechanisms for ordering the magnetic field distributions within CMZ clouds.
The FIREPLACE observations presented in this work have the potential to enhance our understanding of the role magnetic fields play in the star formation process in the complex CMZ. Furthermore, the magnetic field distribution derived from these observations will improve our ability to develop a unified picture of the CMZ-wide magnetic field throughout the GC region.
Appendix A.1 Guide to Individual Fields and CAL Files
We tabulate the fields and calibration files used for the DR2 FIREPLACE data set in Table A.2.1. The files in Table A.2.1 contain calibrated data with pointing corrections applied. Field numbers refer to the locations in Figure 2. “CCW” and “CW” refer to the rotation direction of the major scan direction relative to perpendicular to the Galactic North (30 as discussed in Section 2.1). For completeness, files from Flight 890 are included in this table, though as discussed in the main paper these files are not used in the DR2 reduction. Horizontal lines are included to separate observations from different flights.
Table A.2.2 tabulates the FIREPLACE DR1 files that are combined with the DR2 files shown in Table A.2.1 to produce the full FIREPLACE data release covering the entire CMZ. The format of Table A.2.2 is identical to that of Table A.2.1 except that orientation (“CCW” or “CW”) is not specified since the orientation of the DR1 observations was more variant, as discussed in Section 2.1.
Appendix A.2 Histograms of the 2022 Polarization Pseudovectors
We produce histograms of the polarization pseudovectors of only the FIREPLACE DR2 fields to determine whether any of the magnetic field enhancements observed in Figure 13 are confined to only the DR1 or DR2 portions of the full FIREPLACE data set.
To do so we sort the Nyquist-sampled polarization angles shown in Figure 3 into 10 bins. The histogram representation of the binned polarization angles for the DR2 observations is shown in Figure A.1. As with the histogram of the full FIREPLACE release, an angle of 0 indicates a magnetic field that is aligned with the Galactic plane, with an angle of 90 indicating a magnetic field that is oriented perpendicular to the Galactic plane.
| File Name | Field | Orientation | Scans |
| F0890_HA_POL_09005471_HAWEHWPE_CAL_014-021.fits | 4 | CCW | 8 |
| F0890_HA_POL_09005472_HAWEHWPE_CAL_006-013.fits | 4 | CW | 8 |
| F0890_HA_POL_09005473_HAWEHWPE_CAL_031-038.fits | 5 | CCW | 8 |
| F0890_HA_POL_09005474_HAWEHWPE_CAL_023-030.fits | 5 | CCW | 8 |
| F0890_HA_POL_09005475_HAWEHWPE_CAL_058-061.fits | 6 | CW | 4 |
| F0890_HA_POL_09005475_HAWEHWPE_CAL_064-067.fits | 6 | CW | 4 |
| F0890_HA_POL_09005476_HAWEHWPE_CAL_068-071.fits | 6 | CCW | 4 |
| F0891_HA_POL_09005476_HAWEHWPE_CAL_004-007.fits | 6 | CCW | 4 |
| F0891_HA_POL_09005477_HAWEHWPE_CAL_008-011.fits | 7 | CW | 4 |
| F0891_HA_POL_09005478_HAWEHWPE_CAL_016-023.fits | 7 | CCW | 8 |
| F0891_HA_POL_09005479_HAWEHWPE_CAL_024-031.fits | 8 | CW | 8 |
| F0893_HA_POL_09005479_HAWEHWPE_CAL_009-016.fits | 8 | CW | 8 |
| F0893_HA_POL_09005480_HAWEHWPE_CAL_017-024.fits | 8 | CCW | 8 |
| F0893_HA_POL_09005481_HAWEHWPE_CAL_029-032.fits | 9 | CW | 4 |
| F0893_HA_POL_09005482_HAWEHWPE_CAL_033-040.fits | 9 | CCW | 8 |
| F0893_HA_POL_09005483_HAWEHWPE_CAL_041-044.fits | 10 | CW | 4 |
| F0895_HA_POL_09005483_HAWEHWPE_CAL_003-006.fits | 10 | CW | 4 |
| F0895_HA_POL_09005484_HAWEHWPE_CAL_007-014.fits | 10 | CCW | 8 |
| F0895_HA_POL_09005485_HAWEHWPE_CAL_015-018.fits | 11 | CW | 4 |
| F0895_HA_POL_09005485_HAWEHWPE_CAL_020-023.fits | 11 | CW | 4 |
| F0895_HA_POL_09005486_HAWEHWPE_CAL_024-031.fits | 11 | CCW | 8 |
| F0895_HA_POL_09005487_HAWEHWPE_CAL_032-039.fits | 12 | CW | 8 |
| F0895_HA_POL_09005488_HAWEHWPE_CAL_040-043.fits | 12 | CCW | 4 |
| F0895_HA_POL_09005488_HAWEHWPE_CAL_047-050.fits | 12 | CCW | 4 |
| F0895_HA_POL_09005489_HAWEHWPE_CAL_051-054.fits | 13 | CW | 4 |
| F0916_HA_POL_09005489_HAWEHWPE_CAL_004-011.fits | 13 | CW | 8 |
| F0916_HA_POL_09005490_HAWEHWPE_CAL_013-020.fits | 13 | CCW | 8 |
| F0916_HA_POL_09005490_HAWEHWPE_CAL_022-025.fits | 13 | CCW | 4 |
| F0916_HA_POL_09005493_HAWEHWPE_CAL_026-033.fits | 15 | CW | 8 |
| F0917_HA_POL_09005494_HAWEHWPE_CAL_002-005.fits | 15 | CCW | 4 |
| F0917_HA_POL_09005497_HAWEHWPE_CAL_007-014.fits | 17 | CW | 8 |
| F0917_HA_POL_09005498_HAWEHWPE_CAL_015-022.fits | 17 | CCW | 8 |
| F0918_HA_POL_09005499_HAWEHWPE_CAL_004-011.fits | 18 | CW | 8 |
| F0918_HA_POL_090054100_HAWEHWPE_CAL_012-019.fits | 18 | CCW | 8 |
| F0918_HA_POL_09005496_HAWEHWPE_CAL_020-027.fits | 16 | CCW | 8 |
| File Name | Field | Scans |
| F0775_HA_POL_09005465_HAWEHWPE_CAL_004-011.fits | 1 | 8 |
| F0775_HA_POL_09005466_HAWEHWPE_CAL_012-019.fits | 1 | 8 |
| F0775_HA_POL_09005466_HAWEHWPE_CAL_028-035.fits | 1 | 8 |
| F0775_HA_POL_09005467_HAWEHWPE_CAL_020-027.fits | 2 | 8 |
| F0775_HA_POL_090054111_HAWEHWPE_CAL_036-039.fits | 2 | 4 |
| F0777_HA_POL_09005468_HAWEHWPE_CAL_005-012.fits | 2 | 8 |
| F0777_HA_POL_09005469_HAWEHWPE_CAL_013-020.fits | 3 | 8 |
| F0777_HA_POL_09005470_HAWEHWPE_CAL_021-028.fits | 3 | 8 |
| F0777_HA_POL_09005471_HAWEHWPE_CAL_029-036.fits | 4 | 8 |
| F0777_HA_POL_09005473_HAWEHWPE_CAL_042-049.fits | 5 | 8 |